# arxiv_2024

### v1

@@ Abstract
Multi-agent learning algorithms have been successful at generating superhuman planning in a wide variety of games but have had little impact on the design of deployed multi-agent planners. A key bottleneck in applying these techniques to multi-agent planning is that they require billions of steps of experience. To enable the study of multi-agent planning at this scale, we present GPUDrive, a GPU-accelerated, multi-agent simulator built on top of the Madrona Game Engine that can generate over a million steps of experience per second. Observation, reward, and dynamics functions are written directly in C++, allowing users to define complex, heterogeneous agent behaviors that are lowered to high-performance CUDA. We show that using GPUDrive we are able to effectively train reinforcement learning agents over many scenes in the Waymo Motion dataset, yielding highly effective goal-reaching agents in minutes for individual scenes and generally capable agents in a few hours. We ship these trained agents as part of the code base at https://github.com/Emerge-Lab/gpudrive.
@@ 1 Introduction
Multi-agent learning has been impactful across a wide range of fully cooperative and zero-sum games [cui2023, sophy2022, stratego2022, go2017, population2018, diplomacy2023]. However, its impact on multi-agent planning for settings that mix humans and robots has been muted. In contrast to the ubiquity of multi-agent learning-based agents in zero-sum games, multi-agent planners for most practical robotic systems are not derived from the output of game-theoretically sound learning algorithms. While it is hard to characterize the space of deployed planners since many of them are proprietary, the majority likely use a mixture of collected data for the prediction of human motion and hand-tuned costs. These are then fed into a robust trajectory optimizer or may be based on imitation learning [bronstein22, lu2023imitation]. This approach has been highly effective in scaling up real-world autonomy but can struggle with reasoning about long-term behavior, contingency planning, and interaction with humans in rare, complex scenarios.
@@
  • We provide a multi-agent, GPU-accelerated, and data-driven simulator that runs at over a million steps per second. Our simulator provides provides a testbed for:
@@
  • We provide gym environments in both torch and jax that can be easily configured with standard open-source multi-agent RL and imitation learning libraries. Additionally, we have open-sourced a basic policy-gradient training loop that can be readily used to develop agents.
  • We release implementations of tuned RL algorithms that can process $20$ million steps of experience per hour on consumer-grade GPUs. These can be used to train $95$% goal-reaching agents across $100$ different scenes in two hours on relatively accessible hardware.
  • We open-source strong driving baseline agents that achieve 97% of their goals on a subset of scenes they have been trained to solve. These are integrated into the simulator so that the simulator comes with default, capable, reactive agents.
@@ 2 Related work
There are various open-source frameworks available that support hardware-accelerated reinforcement learning environments. These environments are generally written directly in an acceleration framework such as Numpy [numpy2020], Jax [jax2018github], or Pytorch [torch2024]. In terms of multi-agent accelerated environments, standard benchmarks include JaxMARL[rutherford2023jaxmarl], Jumanji [bonnet2024jumanji], and VMAS [vmas2022] which primarily feature fully cooperative or fully competitive tasks. Each benchmark requires the design of custom accelerated structures per environment. In contrast, GPUDrive focuses on a mixed motive setting and is built atop Madrona, an extensible ECS-based framework in C++, enabling GPU acceleration and parallelization across environments [shacklett2023extensible]. Madrona comes with vectorization of key components of embodied simulation such as collision checking and sensors such as LIDAR. GPUDrive can support hundreds of controllable agents in more than 100,000 distinct scenarios, offering a distinct generalization challenge and scale relative to existing benchmarks. Moreover, GPUDrive includes a large dataset of human demonstrations, enabling imitation learning, inverse RL, and combined IL-RL approaches.
@@
Table 1 shows an overview of current simulators used in autonomous driving research. The purpose of GPUDrive is to facilitate the systematic study of behavioral, coordination, and control aspects of autonomous driving and multi-agent learning more broadly. As such, visual complexity is reduced, which differs from several existing simulators, which (partially) focus on perception challenges in driving [dosovitskiy2017carla, cai2020summit]. Driving simulators close to GPUDrive in terms of either features or speed include MetaDrive [li2022metadrive], nuPlan [caesar2021nuplan], Nocturne [vinitsky2022nocturne], and Waymax [gulino2024waymax] which all utilize real-world data. Unlike MetaDrive and nuPlan, our simulator is GPU-accelerated. Like GPUDrive, Waymax is a JAX-based GPU-accelerated simulator that achieves high throughput through JIT compilation and efficient use of accelerators. With respect to Waymax, our simulator supports a wider range of possible sensor modalities (Section  3.2) including LIDAR and human-like views, and can scale to nearly thirty times more worlds (Section  4.1, and comes with performant reinforcement learning baselines.
@@
Existing simulators often feature baseline agents for interaction, such as low-dimensional car following models that describe vehicle dynamics through a limited set of variables or parameters [kreutz2021analysis, kesting2007general, treiber2000congested]. Rule-based agents exhibit predetermined behaviors, like car-following agents [gulino2024waymax, caesar2021nuplan, SUMO2018, casas2010traffic] such as the IDM model, or parameterized behavior agents like CARLA’s TrafficManager [dosovitskiy2017carla]. Some simulators offer recorded human driving logs for interaction through replaying the human driving logs [lu2023imitation, vinitsky2022nocturne, gulino2024waymax, caesar2021nuplan]. Additionally, certain simulators provide learning-based agents, leveraging reinforcement learning techniques [li2022metadrive]. In our simulator, we provide both human driving logs and high-performing reinforcement learning agents.
@@
TORCS [wymann2000torcs] -
GTA V [martinez2017beyond] -
CARLA [dosovitskiy2017carla] Waypoints
Highway-env [leurent2018environment] -
Sim4CV [muller2018sim4cv] Directions
SUMMIT [cai2020summit] ✓  ($\geq 400$) -
MACAD [palanisamy2020multi] Goal point
SMARTS [zhou2021smarts] Waypoints
MADRaS [santara2021madras] ✓  ($\geq 10$) Goal point
DriverGym [kothari2021drivergym] -
VISTA [amini2022vista] -
nuPlan [caesar2021nuplan] Waypoints
Nocturne [vinitsky2022nocturne] ✓  ($\geq 128$) Goal point
MetaDrive [li2022metadrive] -
InterSim [sun2022intersim] Goal point
TorchDriveSim [scibior2021imagining] -
BITS [xu2023bits] Goal point
Waymax [gulino2024waymax] ✓  ($\geq 128$) Waypoints
@@ 3.1 Simulation Engine
Learning to safely navigate complex scenarios in a multi-agent setting requires generating many billions of environment samples. To feed sample-hungry learning algorithms, GPUDrive is built on top of Madrona [shacklett2023extensible], an Entity-Component-State system designed for high-throughput reinforcement learning environments. In the Madrona framework, multiple independent worlds (each containing an independent number of agentsIn the Waymo Open Motion Dataset, an agent constitutes a vehicle, cyclist, or pedestrian. ) are executed in parallel on accelerators via a shared engine.
However, driving simulation offers a particular set of challenges that require several technical choices. First, road objects, such as road edges and lane lines are frequently represented as polylines (i.e. connected sets of points). These polylines can consist of hundreds of points as they are sampled at every $0.1$ meters, leading to even small maps having upwards of tens of thousands of points. This can blow up the memory requirements of each world as well as lead to significant redundancy in agent observations. Second, the large numbers of agents and road objects can make collision checking a throughput bottleneck. Finally, there is immense variability in the number of agents and road objects in a particular scene. Each world allocates memory to data structures that track its state and accelerate simulation code. Though independent, each world incurs a memory footprint proportional to the maximum number of agents across all worlds. In this way, the performance of GPUDrive is sensitive to the variation in agent counts across all the worlds in a batch.
These challenges are partially resolved via the following mechanisms. First, a primary acceleration data structure leveraged by GPUDrive is a Bounding Volume Hierarchy (BVH). The BVH keeps track of all physics entities and is used to easily exclude candidate pairs for collisions. This allows us to then run a reduced-size collision check on potential collision candidate pairs. The use of a BVH avoids invoking a collision check that would otherwise always be quadratic in the number of agents in a world. Secondly, we observed that a lot of the lines in the geometry of the roads are straight. This allows us to omit many intermediate points while only suffering a minor hit in the quality of the curves. We apply a polyline decimation algorithm (Viswalingham-Whyatt Algorithm) [Visvalingam] to approximate straight lines and filter out low-importance points in the polylines. With this modification, we can reduce the number of points by 10-15 times and significantly improve the step times while decreasing memory usage.
@@ 3.2 Simulator features
Figure 1: Visualization of different observation spaces available in GPUDrive. The top scene is an example scenario from the dataset, rendered from the ego-centric perspective of the red vehicle. Grey cars are parked cars while white cars are other controlled agents. From left to right: the Radius Obs returns all objects within 100 meters, the LIDAR observation with 3000 rays spread around 360 degrees, and a view cone consisting of 3000 rays emanating in a 120-degree view cone.
@@
GPUDrive represents its map as a series of polylines and does not require a connectivity map of the lanes. As such, it can be made compatible with most driving datasets given the pre-processing of the roads into the polyline format. Currently, GPUDrive supports the Waymo Open Motion Dataset (WOMD)  [ettinger2021] which is available under a non-commercial license. The WOMD consists of a set of over 100,000 multi-agent traffic scenarios, each of which contains the following key elements: 1) Road map - the layout and structure of a road, such as a highway or parking garage. 2) Expert human driving demonstrations. 3) Road objects, such as stop signs and crosswalks. Figure 1 depicts an example of an intersection traffic scenario as rendered in GPUDrive.
@@
GPUDrive supports a variety of observation spaces intended to enable heterogeneous types of agents. Fig. 1 depicts the three types of supported state spaces. The first mode is somewhat unphysical in which all agents and road objects within a fixed radius are observable to the agent. This mode is intended primarily for debugging and quick testing, enabling a user to minimize the amount of partial observability in the environment. The other two modes are based on a GPU-accelerated LIDAR scan, representing what an autonomous vehicle would be able to see and what a human would likely be able to see respectively. Both modes are based on casting LIDAR rays; to model human vision we simply restrict the LIDAR rays to emanate in a smaller, controllable-sized cone that can be rotated through an action corresponding to head rotation. Note that since all objects are represented as bounding boxes of fixed height, the LIDAR observations are over-conservative as humans while LIDAR scans in reality are usually able to see over the hoods of cars.
@@
Agents are stepped using a standard Ackermann bicycle model with actions corresponding to steering and acceleration. This model enables the dynamics of objects to be affected by their length, creating different dynamics for small cars vs. trucks. However, this model is not fully invertible which can make it challenging to use this model for imitation learning. To enable full invertibility for imitation learning, we also support the simplified bicycle model, taken from Waymax [gulino2024waymax], which is a double-integrator in the position and velocity and updates its yaw as:
@@
All agents are given a target goal to reach; this goal is selected by taking the last point observed in the vehicle’s logged trajectory. A goal is reached when agents are within some configurable distance $\delta$ of the goal. By default, agents in GPUDrive receive a reward of $1$ for achieving their goal and otherwise receive a reward of 00. There are additional configurable collision penalties or other rewards based on agent-vehicle distances or agent-road distances though these are not used in this work.
@@
We use reinforcement learning to train a set of agents that reach their goals 95 % of the time on a subset of $500$ training scenes. While this number is far below the capability of human drivers, these agents are reactive in a distinct fashion from parametrized driver models in other simulators. In particular, many logged-data simulators construct reactivity by having the driver follow along its logged trajectory but decelerate if an agent passes in front of it. In contrast, these agents can maneuver and negotiate without remaining constrained to a logged trajectory. These trained agents are extremely aggressive about reaching their goals and can be used as an out-of-distribution test for proposed driving agents. The training procedure and more details can be found in Section  4.2.
@@
  • Absence of a map. The current version of the simulator does not have a well-defined notion of lanes or a higher-level road map which makes it challenging for algorithmic approaches that require maps. The absence of this feature also makes it challenging to define rewards such as "stay lane-centered."
@@
  • Unsolvable goals. Due to mislabelling in the Waymo dataset, some agent goals (roughly 2%) are unreachable. For these agents, we default them to simply replaying their logged trajectory and do not treat them as agents.
  • Variance in controllable agents per scenario. In the majority of scenes, there are approximately 8-10 agents and an average of 50 parked cars. Additionally, the dataset is gathered from the sensors of an autonomous vehicle, leading to some agents having their initial states recorded only after the first time-step of the simulator. These agents are not included, as incorporating them would necessitate "teleporting" them into the scene, potentially leading to unavoidable collisions with agents deviating from their logged trajectories.

### v2

# GPUDrive: Data-driven, multi-agent driving simulation at 1 million FPS

@@ Abstract
Multi-agent learning algorithms have been successful at generating superhuman planning in various games but have had limited impact on the design of deployed multi-agent planners. A key bottleneck in applying these techniques to multi-agent planning is that they require billions of steps of experience. To enable the study of multi-agent planning at scale, we present GPUDrive, a GPU-accelerated, multi-agent simulator built on top of the Madrona Game Engine that can generate over a million simulation steps per second. Observation, reward, and dynamics functions are written directly in C++, allowing users to define complex, heterogeneous agent behaviors that are lowered to high-performance CUDA. We show that using GPUDrive we can effectively train reinforcement learning agents over many scenes in the Waymo Open Motion Dataset, yielding highly effective goal-reaching agents in minutes for individual scenes and enabling agents to navigate thousands of scenarios within hours. The code base with pre-trained agents is available at https://github.com/Emerge-Lab/gpudrive.
@@ 1 Introduction
Multi-agent learning has been impactful across a wide range of fully cooperative and zero-sum games (Cui et al., 2023; Wurman et al., 2022; Pérolat et al., 2022; Silver et al., 2017; Jaderberg et al., 2018; Bakhtin et al., 2023). However, its impact on multi-agent planning for settings that mix humans and robots has been muted. In contrast to the ubiquity of multi-agent learning-based agents in zero-sum games, multi-agent planners for most practical robotic systems are not derived from the output of game-theoretically sound learning algorithms. While it is hard to characterize the space of deployed planners since many of them are proprietary, the majority likely use a mixture of collected data for the prediction of human motion and hand-tuned costs. These are then fed into a cost-based trajectory optimizer or into a planner based on imitation learning (Bronstein et al., 2022; Lu et al., 2023). This approach has been highly effective in scaling up real-world autonomy but can struggle with reasoning about long-term behavior, contingency planning, and interaction with humans in rare, complex scenarios.
@@
  • We provide a multi-agent, GPU-accelerated, and data-driven simulator that runs at over a million steps per second (Section 4.1). Our simulator provides a testbed for:
@@
  • We provide gymnasium (Towers et al., 2024) environments in both torch and jax that can be easily configured with standard open-source multi-agent RL and imitation learning libraries. Additionally, we have open-sourced a basic policy-gradient training loop that can be readily used to develop agents.
  • We release implementations of tuned RL algorithms that can process $30$ million steps of experience per hour on consumer-grade GPUs. These can be used to train $95$% goal-reaching agents across 1000 different multi-agent scenarios in 15 hours on relatively accessible hardware.
  • We open-source these strong driving baseline agents that achieve 95% of their goals on a subset of scenes. These are integrated into the simulator so that the simulator comes with default, reactive agents.
@@ 2 Related work
There are various open-source frameworks available that support hardware-accelerated reinforcement learning environments. These environments are generally written directly in an acceleration framework such as Numpy (Harris et al., 2020), Jax (Bradbury et al., 2018), or Pytorch (Ansel et al., 2024). In terms of multi-agent accelerated environments, standard benchmarks include JaxMARL(Rutherford et al., 2023), Jumanji (Bonnet et al., 2024), and VMAS (Bettini et al., 2022) which primarily feature fully cooperative or fully competitive tasks. Each benchmark requires the design of custom accelerated structures per environment. In contrast, GPUDrive focuses on a mixed motive setting and is built atop Madrona, an extensible ECS-based framework in C++, enabling GPU acceleration and parallelization across environments (Shacklett et al., 2023). Madrona comes with vectorization of key components of embodied simulation such as collision checking and sensors such as LIDAR. GPUDrive can support hundreds of controllable agents in more than 100,000 distinct scenarios, offering a distinct generalization challenge and scale relative to existing benchmarks. Moreover, GPUDrive includes a large dataset of human demonstrations, enabling imitation learning, inverse RL, and combined IL-RL approaches.
@@
Table 1 shows an overview of current simulators used in autonomous driving research. The purpose of GPUDrive is to facilitate the systematic study of behavioral, coordination, and control aspects of autonomous driving and multi-agent learning more broadly. As such, visual complexity is reduced, which differs from several existing simulators, which (partially) focus on perception challenges in driving (Dosovitskiy et al., 2017; Cai et al., 2020). Driving simulators close to GPUDrive in terms of either features or speed include MetaDrive (Li et al., 2022), nuPlan (Caesar et al., 2021), Nocturne (Vinitsky et al., 2022), and Waymax (Gulino et al., 2024) which all utilize real-world data. Unlike MetaDrive and nuPlan, our simulator is GPU-accelerated. Like GPUDrive, Waymax is a JAX-based GPU-accelerated simulator that achieves high throughput through JIT compilation and efficient use of accelerators. With respect to Waymax, our simulator supports a wider range of possible sensor modalities (Section  3.2) including LIDAR and human-like views, can scale to nearly thirty times more worlds (see Section  4.1), and comes with performant reinforcement learning baselines. However, it does not currently come with reactive IDM agents like Waymax though it does come with pre-trained simulated agents based on RL policies.
@@
Existing simulators often feature baseline agents for interaction, such as low-dimensional car following models that describe vehicle dynamics through a limited set of variables or parameters (Kreutz & Eggert, 2021; Kesting et al., 2007; Treiber et al., 2000). Rule-based agents exhibit predetermined behaviors, like car-following agents (Gulino et al., 2024; Caesar et al., 2021; Lopez et al., 2018; Casas et al., 2010) such as the IDM model, or parameterized behavior agents like CARLA’s TrafficManager (Dosovitskiy et al., 2017). Some simulators offer recorded human driving logs for interaction through replaying the human driving logs (Lu et al., 2023; Vinitsky et al., 2022; Gulino et al., 2024; Caesar et al., 2021). Additionally, certain simulators provide learning-based agents, leveraging reinforcement learning techniques (Li et al., 2022). In GPUDrive, we provide both human driving logs and high-performing reinforcement learning agents.
@@
TORCS (Wymann et al., 2000) -
GTA V (Martinez et al., 2017) -
CARLA (Dosovitskiy et al., 2017) Waypoints
Highway-env (Leurent et al., 2018) -
Sim4CV (Müller et al., 2018) Directions
SUMMIT (Cai et al., 2020) ✓  ($\geq 400$) -
MACAD (Palanisamy, 2020) Goal point
SMARTS (Zhou et al., 2021) Waypoints
MADRaS (Santara et al., 2021) ✓  ($\geq 10$) Goal point
DriverGym (Kothari et al., 2021) -
VISTA (Amini et al., 2022) -
nuPlan (Caesar et al., 2021) Waypoints
Nocturne (Vinitsky et al., 2022) ✓  ($\geq 128$) Goal point
MetaDrive (Li et al., 2022) -
InterSim (Sun et al., 2022) Goal point
TorchDriveSim (Ścibior et al., 2021) -
BITS (Xu et al., 2023) Goal point
Waymax (Gulino et al., 2024) ✓  ($\geq 128$) Waypoints
@@ 3.1 Simulation Engine
Learning to safely navigate complex scenarios in a multi-agent setting requires generating many billions of environment samples. To feed sample-hungry learning algorithms, GPUDrive is built on top of Madrona (Shacklett et al., 2023), an Entity-Component-State system designed for high-throughput reinforcement learning environments. In the Madrona framework, multiple independent worlds (each containing an independent number of agentsIn the Waymo Open Motion Dataset, an agent constitutes a vehicle, cyclist, or pedestrian. ) are executed in parallel on an accelerator via a shared engine.
However, the simulation of driving poses a unique set of challenges which require careful technical design choices to solve. First, road objects, such as road edges and lane lines are frequently represented as polylines (i.e. connected sets of points). These polylines can consist of hundreds of points as they are sampled at every $0.1$ meters, leading to even small maps having upwards of tens of thousands of points. This can blow up the memory requirements of each world as well as lead to significant redundancy in agent observations. Second, the large numbers of agents and road objects can make collision checking a throughput bottleneck. Finally, there is immense variability in the number of agents and road objects in a particular scene. Each world allocates memory to data structures that track its state and accelerate simulation code. Though independent, each world incurs a memory footprint proportional to the maximum number of agents across all worlds. In this way, the performance of GPUDrive is sensitive to the variation in agent counts across all the worlds in a batch.
These challenges are partially resolved via the following mechanisms. First, a primary acceleration data structure leveraged by GPUDrive is a Bounding Volume Hierarchy (BVH). The BVH keeps track of all physics entities and is used to easily exclude candidate pairs for collisions. This allows us to then run a reduced-size collision check on potential collision candidate pairs. The use of a BVH avoids invoking a collision check that would otherwise always be quadratic in the number of agents in a world. Secondly, we observed that a lot of the lines in the geometry of the roads are straight. This allows us to omit many intermediate points while only suffering a minor hit in the quality of the curves. We apply a polyline decimation algorithm (Visvalingham-Whyatt Algorithm) (Visvalingam & Whyatt, 1993) to approximate straight lines and filter out low-importance points in the polylines. With this modification, we can reduce the number of points by 10-15 times and significantly improve the step times while decreasing memory usage. Finally, rather than allocate memory for the maximum number of agents in a scene (as is likely necessary in frameworks like Jax), we only allocate memory equal to the actual number of instantiated agents.
@@ 3.2 Simulator features
Figure 1: Four example scenarios from the Waymo Open Motion Dataset rendered in GPUDrive. The blue circles indicate vehicle destinations.
@@
GPUDrive represents its map as a series of polylines and does not require a connectivity map of the lanes. As such, it can be made compatible with most driving datasets given the pre-processing of the roads into the polyline format. Currently, GPUDrive supports the Waymo Open Motion Dataset (WOMD)  (Ettinger et al., 2021) which is available under a non-commercial license. We show four representative example scenarios from the WOMD in Figure 1. The WOMD consists of a set of over 100,000 multi-agent traffic scenarios, each of which contains the following key elements: 1) Road map - the layout and structure of a road, such as a highway or parking garage. 2) Logged human trajectories from vehicles, cyclists, and pedestrians. 3) Road objects, such as stop signs and crosswalks. Figure 2 depicts an example of an intersection traffic scenario as rendered in GPUDrive.
@@
GPUDrive supports a variety of observation spaces intended to enable heterogeneous types of agents. Fig. 2 depicts the three types of supported state spaces. The first mode is somewhat unphysical in which all agents and road objects within a fixed radius are observable to the agent. This mode is intended primarily for debugging and quick testing, enabling a user to minimize the amount of partial observability in the environment. The other two modes are based on a GPU-accelerated LIDAR scan, representing what an autonomous vehicle would be able to see and what a human would likely be able to see respectively. Both modes are based on casting LIDAR rays; to model human vision we simply restrict the LIDAR rays to emanate in a smaller, controllable-sized cone that can be rotated through an action corresponding to head rotation. Note that since all objects are represented as bounding boxes of fixed height, the LIDAR observations are over-conservative as in reality it is frequently possible to see over the hoods of cars as their height is lower than the body of the rest of the car.
Figure 2: Visualization of different observation spaces available in GPUDrive. The top scene is an example scenario from the dataset, rendered from the ego-centric perspective of the red vehicle. Grey cars are parked cars while white cars are other controlled agents. From left to right: the Radius Filter returns all objects within 100 meters, the LIDAR observation with 3000 rays spread around 360 degrees, and a view cone consisting of 3000 rays emanating in a 120-degree view cone.
@@
By default, agents are stepped using a standard Ackermann bicycle model (Details in Appendix B) with actions corresponding to steering and acceleration. This model enables the dynamics of objects to be affected by their length, creating different dynamics for small cars as opposed to large cars like trucks. However, this model is not fully invertible which can make it challenging to use as a model for imitation learning. To enable full invertibility for imitation learning, we also support the simplified bicycle model, taken from Waymax (Gulino et al., 2024), which is a double-integrator in the position and velocity and updates its yaw as:
@@
All agents are given a target goal to reach; this goal is selected by taking the last recorded position in the vehicle’s logged trajectory. A goal is reached when agents are within some configurable distance $\delta$ of the goal. By default, agents in GPUDrive receive a reward of $1$ for achieving their goal and otherwise receive a reward of 00. There are additional configurable collision penalties or other rewards based on agent-vehicle distances or agent-road distances though these are not used in the experiments reported in this work.
Termination conditions.
We terminate an agent’s episode when they achieve their goal position or collide. As it is not clear where an agent should go next after it reaches its goal, we simply remove it from the scene afterwards (we note that an alternative might be to generate a new goal for the agent to drive to). Car and cyclist agents are considered to be in collision when they drive through a road edge while pedestrian agents are allowed to cross road edges.
Environment Interface.
As GPUDrive is implemented in C++, we provide a Pythonic interface through nanobind (Jakob, 2022). We create environments for both torch and jax that conform to the Gymnasium API (Towers et al., 2024) so users can use the simulator entirely through Python if they prefer.
@@
We use reinforcement learning to train a set of agents that reach their goals 95 % of the time on a subset of $1000$ training scenes. While this number is far below the capability of human drivers, these agents are reactive in a distinct fashion from parametrized driver models in other simulators. In particular, many logged-data simulators construct reactivity by having the driver follow along its logged trajectory but decelerate if an agent passes in front of it. In contrast, these agents can maneuver and negotiate without remaining constrained to a logged trajectory. These trained agents are extremely aggressive about reaching their goals and can be used as an out-of-distribution test for proposed driving agents. The training procedure and more details can be found in Section  4.2.
@@
  • Absence of a map. The current version of the simulator does not have a well-defined notion of lanes or a higher-level road map which makes it challenging for algorithmic approaches that require maps. The absence of this feature also makes it challenging to define rewards such as ”stay lane-centered.”
@@
  • Unsolvable goals. Due to incorrect labels of some road points in the Waymo dataset, such as an exit to a parking lot being labelled as an impassable road edge, some agent goals (roughly 2%) are unreachable. For these agents, we default them to simply replaying their logged trajectory and do not treat them as agents.
  • Variance in controllable agents per scenario. In most scenes, a significant fraction (25-75%) of the agents are not controllable because they do not move (e.g. parked cars). This is because the dataset is collected from an autonomous vehicle’s sensors, so some agents’ initial states are recorded after the first time-step of the simulator. These agents are not included to avoid collisions with other agents.